# Grounding Multimodal LLMs to Embodied Agents that Ask for Help with Reinforcement Learning

### v2

@@ 5.4 Analysis
Figure 4: Success Rate vs. Number of Required and Asked Questions. Evaluation performance of SFT and RL trained policies vs. number of required and total questions asked by the agent.
Quantitative analysis of agent behavior. In Fig. 4 (a.), we show the performance of MLLM policies trained using RL and SFT vs. minimum required questions to solve tasks on Unseen Tasks split. These results show, as the number of required questions for a novel task increases the performance of both methods drops. In Fig. 4 (b.), we show the performance of these policies vs. number of questions asked by the agents. This results shows, as the number of questions asked by agents increase the performance starts dropping (from $0-4$ questions for SFT and $1-10$ questions for RL), however both methods there’s a inflection point after which the performance increases (between $4-5$ for SFT and $10-15$ for RL). This suggests as these agents ask a lot more questions than minimum required for a task they are able to resolve ambiguity. We also observe, agents trained with SFT tend to be more conservative asking fewer questions compared to RL-trained counterparts that could explain the difference in performance as shown in Tab. 2 (row 5 vs. 6).
Figure 5: Qualitative Example. Successful trajectory of our method on an evaluation episode from Unseen Tasks split.
Qualitative examples. In Fig. 4, we show a success example of our method on Unseen Tasks split. The agent first explores until it finds all $4$ instances of the cup. Then it asks a sequence of questions. Finally, based on responses from the user it successfully brings the green cup user wants to the light table. For more qualitative examples, see Appendix E.
@@ Appendix A Dataset
In Fig. 6, we show additional examples of different types of ambiguities in our Ask-to-Act dataset. Specifically we show 3 additional examples for:
  • Attribute Recognition: The agent must reason about appearance-based attributes such as color and object category to disambiguate the object user is looking for (e.g., a red bowl and white bowl as shown in Fig. 6 (a)).
  • Object Size : The agent must distinguish between objects of different sizes but similar appearance and location to disambiguate the object user is looking for (e.g., a large red bowl vs. a small red bowl as shown in Fig. 6 (b)) .
  • Attribute/Spatial Reasoning and Object Size: A combination of attribute and spatial reasoning with object size, where the agent must account for both appearance or location and size to effectively disambiguate the target object the user is looking for. For the example shown in Fig. 6 (c), the agent needs to reason about different colored shoes (blue, white, and orange shoe) which are located on the blue cabinet and the sofa and some of those have different size e.g. large blue shoe on blue cabinet vs. small blue shoe on sofa and large white shoe on sofa vs. small white shoe on blue cabinet.
Object Categories. We present the full list of object categories used for Ask-to-Act dataset in Tab. 4.
@@
Appendix B Additional Experiments
Unseen Scenes Unseen Tasks
[…]
10) Random + GT Planner $37.3$- -- $23.7$- --
11) Ours w/o ask questions $33.7$ $25.7$ $0.0$ $22.4$ $9.2$ $0.0$
12) Ours $89.8$ $63.2$ $2.6$ $65.2$ $32.4$ $2.5$
Table 5: Ablations. Comparison of RL training with and without ask question actions on Interact-and-Ask benchmark.
Can agents succeed without asking for help? In Tab. 5, we show a experiment where we ask whether it is possible to train an effective MLLM policy using RL that generalizes well to unseen scenarios in Ask-to-Act task without asking a single question. We first set up a random baseline for Ask-to-Act, we use privileged information from the simulator to fetch the list of objects of target object category for each episode and randomly choose one of them as the target object. Next, we use the ground truth single object rearrangement planner we wrote in Habitat 3.0 [30] to place the randomly sampled target object on the target receptacle. Results of evaluating this random baseline with ground truth planner is shown in row 1 of  Tab. 5. Next, we use the same policy architecture, and reward as described in our approach in Sec. 4 and modify the action space of the policy to disable ‘ask_question‘ question and train it with RL i.e. the agent cannot ask the user (simulated by the LLM in our task) for clarification questions. As shown by results in row 2 of  Tab. 5, the MLLM policy trained to solve our task without asking questions achieves performance close to a random baseline (row 1 vs. row 2). In contrast, a policy trained using our LLM generated rewards with RL (row 3) achieves significantly better performance on both Unseen Scenes and Unseen Tasks evaluation splits.
Appendix C Training Details
@@
Training Details. To train our policy using reinforcement learning (RL) we use PPO with Generalized Advantage Estimation (GAE) [35]. We use a discount factor $\gamma$ of $0.99$ and set GAE parameter $\tau$ to 0.95. We do not use normalized advantages. To parallelize training, we use DD-PPO [47], an adaptation of PPO [36] for distributed training, with $24$ environments each on $8$ A40 GPUs for $50$ million steps . Each worker collects $24$ frames of experience from $24$ environments parallely and then performs $2$ epochs of PPO update with $2$ mini batches in each epoch. Tab. 6 details the default hyperparameters used in all of our training runs.
@@
To compute action probabilities, one naive approach is to take a sum over all token-level probabilities. When training MLLMs with action-space with variable number of tokens across actions one issue in  Eq. 3, is that actions with larger number of tokens tend to have lower token-level probabilities, even though they might be more reasonable to take at some point in the environment. This issue happens because the probability of each token $P(u_{t}^{i}|\cdot)$ is always less than $1$. This could be problematic in RL optimization especially for the case when the agent needs to output a natural language question for Ask-to-Act task, simply because the questions tend to have more tokens than pre-defined skills represented in language. To remedy this issue, we use action length normalization technique ınormalize the token-level probabilities of the actions with the number of action tokens, which can be defined as:
@@
We find action-length normalization helps stabilize training when working with our variable length action-space when using constrained grammar decoding [29] for RL training.
Appendix D LLM Reward Generation
@@
4. ask("question")
You can only ask the following 3 types of yes/no questions about the object in question:
1. Is it on the <receptacle>?
2. Is it the <object_instance>?
3. Is it the <object_size> one?
@@
Table 7: LLM prompt used for generating optimal question sequences and subgoals used to generate rewards and build the answering module.
Tab. 7 shows the prompt used for generating optimal answer sequences and subgoals required to generate the reward function described in Eq. 1 in Sec. 4.2. The same outputs are also used to build a answering module (described in Sec. 3) used during training and evaluation.
$r_{t}=$ $10\cdot\mathds{1}_{\text{success}}+2.5\cdot\mathds{1}_{\text{subgoal}}+r_{3}\cdot\mathds{1}_{\text{useful\_question}}$ (5)
$-0.05\cdot\mathds{1}_{\text{exceed\_budget}}-0.01,$
Reward Details. Eq. 5 shows the reward function we use with coefficients for each term. In this reward, $\mathds{1}_{\text{success}}$ indicates whether the task was successfully completed and $\mathds{1}_{\text{subgoal}}$ (generated by the LLM) indicates if the agent completed any subgoal required to complete the overall task. For example, for a task “Bring me the cup and place it on the coffee table”, the agent needs to first search for all cups, pick the correct cup, then navigate to the coffee table, and finally place it. Similarly, $\mathds{1}_{\text{useful\_question}}$ (generated by the LLM) indicates if a question asked by the agent is valid and helps make progress towards disambiguating the target object or not. Consider the example of fetching the cup, if the environment has $4$ cups on a table and the agent asks “Is it the red cup?” and if user responds ‘no’, then the agent should ask ‘Is it the blue cup?’ or ‘Is it the yellow cup?’ to find the target object instead of asking ‘Is it on the table?’ to make progress towards solving the task. Additionally, for each question the agent asks in an episode the reward $r_{3}$ is a value normalized to sum up to $1$ based on number of questions required to solve the task. For example, if the task requires $3$ questions then the agent will get $r_{3}=0.33$ for each question it gets correct from relevant question subset until agent asks $3$ relevant questions and 00 afterwards. When training policies under a budget of questions we penalize the agent for every question that exceeds pre-specified budget given by $\mathds{1}_{\text{exceed\_budget}}$. By default the budget is set to minimum number of required questions $K$ to solve the task.
Appendix E Qualitative Examples
We present additional qualitative examples of evaluating our method on Unseen Tasks evaluation split of Ask-to-Act dataset in Fig. 7.
Figure 7: Qualitative Example. Successful trajectories of our method on $2$ evaluation episodes from Unseen Tasks split.
Appendix F Baseline Details
(a.) Fully Observable Text WorldGraph + ReAct (Zero-shot). In this baseline, we provide an LLM (GPT4o in our case) with a fully observable text-based world graph describing the environment, including receptacles, objects, and their locations (e.g. “The apple is on the coffee table.). The prompt used for this baseline is shown in Tab. 8.
(b.) Fully Observable Text WorldGraph + ReAct (Few-shot). This baseline extends (a) by providing the LLM with a few in-context examples that demonstrate task planning and ambiguity resolution strategies in the Ask-to-Act task. By leveraging demonstrations, this approach assesses whether in-context learning improves LLMs task planning and ambiguity reasoning under full observability. The prompt used for this baseline is shown in Tab. 9.
(c.) Partially Observable Text WorldGraph + ReAct (Few-shot). While baselines (a) and (b) assume privileged access to a fully observable world graph, constructing such representations in real-world settings is often infeasible. This baseline relaxes that assumption by providing the LLM with a partially observable text-based world graph. At the start of an episode, the agent lacks full knowledge of object locations and must actively explore to gather necessary information. The prompt used for this baseline is shown in Tab. 10.
(d.) Vision GPT4o + SoM + ReAct. Building an error-free text representation of real-world environments is challenging. This baseline evaluates whether existing MLLMs can solve the Ask-to-Act task using egocentric visual observations. At each timestep, the MLLM receives the robots visual input along with an skill library for executing actions. To enhance grounding, we label visual observations using Set-of-Marks (SoM) [52] and maintain memory by providing GPT4o with a textual history of past observations and actions. The prompt used for this baseline is shown in Tab. 11, in addition to the prompt this baseline takes in visual observations from robotos egocentric camera augmented with SoM [52].
Appendix G Limitations
In this work, the Ask-to-Act tasks are purposefully constrained in terms of the types of questions that can be asked so that we could study this problem in a controlled manner. For example, this made the evaluation of questions feasible in our setting, such that we can focus on evaluating whether it is possible to train an embodied agent that can interleave task planning with multi-round question-answering with the user. We show in this paper that it is indeed possible, and we leveraged the structure to disentangle the types of errors that can occur. An important future work is to study a more realistic setup of Ask-to-Act with a much more complex set of ambiguous tasks and an open-ended question-answer setup. There are a range of open research questions required to do this, including the use (and potentially training/distillation) of efficient LLM-as-reward models that can be used within an RL pipeline. Further study is also needed to determine exactly how accurate or reliable the rewards must be to train effective models. As it is anticipated that LLM-generated rewards will not be as accurate in less constrained settings. Additionally, using an LLM in an open-ended question answering setting while collecting embodied experience require solving non-trivial infrastructure challenges, such as running a LLM within the training loop for both question-answering and reward modeling with a simulator for tens of millions of steps. By constraining the types of questions an agent can ask we could make RL training tractable to study the effectiveness of LLM-generated rewards for Ask-to-Act in a controller manner. These aspects are limitations of our current work but exciting research directions for future work.
@@ Appendix G Limitations
4. ask("question")
You can only ask the following 5 types of yes/no questions about the object in question:
1. Is it on the <receptacle>?
2. Is it the <object_instance>?
3. Is it the <object_size> one?
After you ask a question wait for the human to reply.
Strictly follow the above format while solving given task.
Instruction: instruction
Receptacles: receptacles
@@ 3. Is it the <object_size> one?
receptacles_with_objects
Observation action history:
prev_observations
Current observation:
Previous Actions: prev_actions
Agent at: current_room
Thought:

### v3

@@ 5.4 Analysis
Quantitative analysis of agent behavior. In Fig.˜3 (c.), we show the performance of MLLM policies trained using RL and SFT vs. minimum required questions to solve tasks on Unseen Tasks split. These results show, as the number of required questions for a novel task increases the performance of both methods drops. In Fig.˜3 (d.), we show the performance of these policies vs. number of questions asked by the agents. This shows, as the number of questions asked by RL agents increase (from 131-3 questions) the task success rate starts increasing. We also observe, agents trained with SFT tend to be more conservative asking fewer questions compared to RL-trained counterparts that could explain the difference in performance as shown in Tab.˜1 (row 5 vs. 6).
Figure 4: Qualitative Example. Successful trajectory on an evaluation episode from Unseen Tasks split.
Qualitative examples. Fig.˜4 shows a successful example of our method on Unseen Tasks split. The agent first explores until it finds all $4$ instances of the cup then it asks a sequence of questions. Finally, based on responses from the user it successfully brings the green cup user wants to the light table. For more qualitative examples, see Appendix˜D.
@@ Appendix A Dataset
In Fig.˜5, we show additional examples of different types of ambiguities in our Ask-to-Act dataset. Specifically we show 3 additional examples for:
  • Attribute Recognition: The agent must reason about appearance-based attributes such as color and object category to disambiguate the object user is looking for (e.g., a red bowl and white bowl as shown in Fig.˜5 (a)).
  • Object Size: The agent must distinguish between objects of different sizes but similar appearance and location to disambiguate the object user is looking for (e.g., a large red bowl vs. a small red bowl as shown in Fig.˜5 (b)) .
  • Attribute/Spatial Reasoning and Object Size: A combination of attribute and spatial reasoning with object size, where the agent must account for both appearance or location and size to effectively disambiguate the target object the user is looking for. For the example shown in Fig.˜5 (c), the agent needs to reason about different colored shoes (blue, white, and orange shoe) which are located on the blue cabinet and the sofa and some of those have different size e.g. large blue shoe on blue cabinet vs. small blue shoe on sofa and large white shoe on sofa vs. small white shoe on blue cabinet.
Object Categories. We present the full list of object categories used for Ask-to-Act dataset in Tab.˜3.
@@
Appendix B Training Details
@@
Training Details. To train our policy using reinforcement learning (RL) we use PPO with Generalized Advantage Estimation (GAE) (Schulman et al., 2016). We use a discount factor $\gamma$ of $0.99$ and set GAE parameter $\tau$ to 0.95. We do not use normalized advantages. To parallelize training, we use DD-PPO (Wijmans et al., 2020), an adaptation of PPO (Schulman et al., 2017) for distributed training, with $24$ environments each on $8$ A40 GPUs for $50$ million steps . Each worker collects $24$ frames of experience from $24$ environments parallely and then performs $2$ epochs of PPO update with $2$ mini batches in each epoch. Tab.˜4 details the default hyperparameters used in all of our training runs.
Per-Step Reward Inference. For per-step reward inference for evaluating questions asked by an agent, we use a vLLM (Kwon et al., 2023) server to host our LLM answer evaluation model using Llama-3.1 8B Instruct (Grattafiori et al., 2024) model. At each rollout step during RL training and policy evaluation, if the agent outputs a ‘ask_question’ action we send a request to the vLLM server and feed the response back to the policy at the next step. In order to achieve a high throughput we run 8 parallel vLLM servers and randomly send requests to different workers from training and evaluation workers.
@@
To compute action probabilities, one naive approach is to take a sum over all token-level probabilities. When training MLLMs with action-space with variable number of tokens across actions one issue in Eq.˜3, is that actions with larger number of tokens tend to have lower token-level probabilities, even though they might be more reasonable to take at some point in the environment. This issue happens because the probability of each token $P(u_{t}^{i}|\cdot)$ is always less than $1$. This could be problematic in RL optimization especially for the case when the agent needs to output a natural language question for Ask-to-Act task, simply because the questions tend to have more tokens than pre-defined skills represented in language. To remedy this issue, we use action length normalization technique normalize the token-level probabilities of the actions with the number of action tokens, which can be defined as:
@@
We find action-length normalization helps stabilize training when working with our variable length action-space when using constrained grammar decoding (Park et al., 2025) for RL training.
Appendix C LLM Reward Generation
@@
4. open(receptacle) // Strictly use when the object needs be to picked from or placed inside an articulated receptacle after navigating to it
5. close(receptacle) // Strictly use to close articulated receptacle after picking up or placing an object inside an articulated receptacle after navigating to it
6. ask(question)
You can ask the following 9 types of questions about the object in question:
1. Is target object on the <receptacle>?
2. Is <object_instance> the target object?
3. Is target object the <object_size> one?
4. Where is the <object_category> located?
5. What color is <object_category>?
6. Can you describe the <object_category>?
7. Is <object_instance> clutter?
8. Are <object_category> clutter?
9. Which <receptacle> to place the <object_instance> on/in?
@@
Table 5: LLM prompt used for generating optimal question sequences used to generate SFT data.
@@
4. open(receptacle) // Strictly use when the object needs be to picked from or placed inside an articulated receptacle after navigating to it
5. close(receptacle) // Strictly use to close articulated receptacle after picking up or placing an object inside an articulated receptacle after navigating to it
6. ask(question)
Your task is to evaluate the questions asked by an embodied agent to identify the target object and resolve ambiguity about placement preference if any. You will be provided with task instruction, objects in environment, current location of objects in the environment, list of target objects and their current location, and desired placement location of each target object. Respond with a concise answer with either object name, receptacle name, object attributes and appearance, or yes/no response followed by a boolean that denotes whether the question is useful towards making progress to resolve ambiguity or not. Here is the example task:
Instruction: Place the bowl in kitchen counter drawer
Receptacles: [light table, chair, sofa, dark table, tv stand, cabinet, sink, left kitchen counter drawer, right kitchen counter drawer, top kitchen counter drawer, bottom kitchen counter drawer, fridge]
@@
coffee table: [blue casserole]
dark table: [red bowl, yellow bowl, black toy]
light table: [yellow dumbbell, blue bowl, red bowl]
Target Object Metadata: {"name": "red bowl", "current location": "on light table", "target location": cabinet}
Questions asked so far: []
Current question: ask("where is the red bowl?")
Answer:
Table 6: LLM prompt used for generating reward for evaluating ‘ask question’ action during RL training.
Tab.˜6 shows the prompt used by the question evaluation LLM used for Ask-to-Act task evaluation and RL training. In addition, we use the prompt 5 for generating optimal answer sequences and subgoals required to generate the SFT data and task subgoals described in Eq.˜1 in Sec.˜4.2.
[…]
Reward Details. Eq.˜5 shows the reward function we use with coefficients for each term. In this reward, $\mathds{1}_{\text{success}}$ indicates whether the task was successfully completed and $\mathds{1}_{\text{subgoal}}$ (generated by the LLM) indicates if the agent completed any subgoal required to complete the overall task. For example, for a task “Bring me the cup and place it on the coffee table”, the agent needs to first search for all cups, pick the correct cup, then navigate to the coffee table, and finally place it. Similarly, $\mathds{1}_{\text{useful\_question}}$ (generated by the LLM) indicates if a question asked by the agent is valid and helps make progress towards disambiguating the target object or not. Consider the example of fetching the cup, if the environment has $4$ cups on a table and the agent asks “where is it the cup?” and if user responds ‘on the coffee table’, then the agent should ask ‘Is it the blue cup?’ or ‘Is it the yellow cup?’ to find the target object instead of asking ‘Is it blue cup in the sink?’ to make progress towards solving the task. For each useful question $\mathds{1}_{\text{useful}}$ the agent asks in an episode it is given a reward $r_{3}$ until total number of questions asked are less than the question budget. For example, if the task requires $3$ questions then the agent will get $r_{3}=0.5$ for each relevant question it asks as evaluated by the LLM and 0 otherwise or if the question budget is exceeded. When training policies under a budget of questions we penalize the agent for every question that exceeds pre-specified budget given by $\mathds{1}_{\text{exceed\_budget}}$. By default the budget is set to minimum number of required questions $K$ to solve the task.
Appendix D Qualitative Examples
We present additional qualitative examples of evaluating our method on Unseen Tasks evaluation split of Ask-to-Act dataset in Fig.˜6.
Figure 6: Qualitative Example. Successful trajectories of our method on 22 evaluation episodes from Unseen Tasks split.
Appendix E Baseline Details